# GEM: Gaussian Embedding Modeling for Out-of-Distribution Detection in GUI Agents

### v2

@@ 5.1 Experiments Setup
[…]
Implementation. Popular GUI Agents wu2024atlas ; qin2025ui ; xu2024aguvis are developed based on Qwen2-VL-7B. To simulate the construction process of a popular GUI Agent, we perform SFT on Qwen2-VL-7B using AITZ train dataset, and then evaluate its OOD detection performance on each sample from eight different OOD datasets and AITZ test dataset (ID). We report the accuracy, precision, recall, and F1 score for each dataset. Detailed descriptions of the experiments is provided in Appendix D.
Baseline. We use two commonly-used types of OOD detection methods as our experimental baselines: embedding-based methods and uncertainty-based methods. The embedding-based methods include the TV score wang2024embedding , last layer embedding, and best layer embedding. The uncertainty-based methods include top-k confidence and output entropy. More detailed descriptions and formal definitions of these baseline methods are provided in Appendix E.
5.2 Main Results
6 Further analysis
In this section, we present several interesting observations based on our proposed method as well as traditional OOD detection methods.
6.1 Generalization Experiment
To demonstrate the generalization ability of GEM, we evaluate its performance across nine different GUI agents or MLLMs. As shown in Table 3, GEM consistently achieves high OOD detection accuracy across five models spanning three platforms, indicating its strong generalization capability.
For the other four models where GEM performs suboptimally, the results remain explainable. Specifically, OS-Atlas-Base-7B and OS-Atlas-Pro-7B were exposed to most of the smartphone datasets used in our experiment during their GUI grounding pretraining phase, and their visual encoders were not frozen, which contributed to their weaker OOD performance on the smartphone platform. Meanwhile, LLaVA-1.5 and BLIP capture visual information at a coarser granularity. Although they still achieve respectable OOD detection accuracy (72.44% to 88.38%) in internalization-OOD scenarios involving the computer and web browser platforms, their performance degrades when both the ID and OOD datasets are from the smartphone domain. This highlights their limitations in handling the complex embedding space specific to GUI agent tasks. This also indicates that extrapolation-OOD scenarios are more challenging for OOD detection than internalization-OOD scenarios.
We also demonstrate the rationality and stability of GEM through ablation studies (see Appendix G).
Table 3: The performance of GEM with different encoder structures.
Smartphone Computer Web browser
Model Accuary F1 score Accuary F1 score Accuary F1 score
@@ 6.1 Generalization Experiment
6.2 Layer-Level OOD Detection Analysis of GUI Agents
We evaluated the effectiveness of representations extracted from each of the twenty-eight layers of Qwen2-VL-7B for OOD detection using an embedding-based approach.
As shown in Figure 5, in most datasets, the AUROC first increases as the layer depth increases, then decreases. However, the AUROC rises again in the last two layers. Interestingly, in 4 out of 8 datasets, the best OOD detection results are achieved at the ninth layer. This might be because, as the layers get deeper, the importance of specific task-related features increases, while the contribution of general visual or textual features decreases. Around the ninth layer, the GUI agent likely finds a balance between these two factors. The increase in AUROC in the final two layers might be because, at this stage, the GUI agent has developed a kind of confidence estimation in its final output, allowing it to better distinguish between OOD and ID samples.
@@ 6.2 Layer-Level OOD Detection Analysis of GUI Agents
7 Conclusion
We present the first systematic analysis of OOD detection for GUI agents and show that traditional OOD detection methods perform suboptimally in this domain. We find that the multimodal inputs of GUI agents exhibit a clustered pattern in the semantic space based on their distance from the centroid. Based on this, we propose GEM, a method that uses input embedding information to fit a GMM to detect OOD samples for GUI agents. Experiments on eight datasets across three platforms demonstrate the superior performance of GEM over all baselines. Generalization experiments on nine different GUI agents or MLLM backbones validate the excellent generalization ability of GEM. We also find that for the GUI agent built with Qwen2-VL-7B, semantic information around the ninth layer achieves the best OOD detection performance.
Appendix A Ethics Statement
@@
Appendix B Expanded Analysis
Figure 6: GUI agent performance across different cluster distances.
In this section, we provide an expanded analysis of content not fully covered in the main text. First, we discuss our limitations, and then we present a detailed analysis of the clustering phenomenon in the input semantic space of the GUI agent.
B.1 Limitations
@@
B.2 The detailed analysis of clustering phenomenon
@@
Appendix C Datasets Details
@@
$\bullet$ AITZ is the first dataset to annotate Chain-of-Action-Thought (CoAT) for GUI agents, which considers descriptions of prior actions and the current screen, the actions that should be taken, and reasoning about the potential consequences of the selected actions. It contains 18,643 screen-action pairs with CoAT annotations.
$\bullet$ AndroidControl is a dataset covering 833 distinct Android applications and containing 15,283 everyday tasks performed within these apps. This dataset is currently the most diverse benchmark for computer-based control, providing a foundation for in-depth model performance analysis.
$\bullet$ OS-Kairos is a dataset containing data from 12 different Chinese and English mobile applications across 12 topic categories. The tasks in this dataset were automatically collected by an agent capability probing framework and later curated by humans. It covers a wide range of scenarios encountered in everyday life.
$\bullet$ Meta-GUI is a dataset containing dialogues and GUI interaction trajectories. It covers six high-frequency daily-life topics across 11 mobile applications.
$\bullet$ Omniact is a dataset consisting of 9.8K pairs of screen images and natural language task descriptions, covering a variety of operating systems and web applications. The dataset includes tasks of varying difficulty levels, ranging from simple single-step actions to complex multi-step operations.
$\bullet$ ScreenSpot is a dataset containing high-resolution real screenshots and expert annotations from various domains. This dataset covers data from multiple platforms, including Mac, Windows, Android, iPhone, and web, making it the most platform-diverse GUI dataset.
Appendix D Main Experiment Details
@@
Evaluation Metrics Details. After performing SFT on Qwen2-VL-7B using the AITZ training set, we fit a distance-based GMM on the same training set following the GEM method. Classification boundaries are defined using three standard deviations per cluster. The accuracy, precision, recall, and F1 scores reported in Table 2 are obtained by treating each dataset as the OOD samples and the AITZ test set as the ID samples. The prompt of GEM and all baselines is shown in Figure 7
@@ E.1 TV score
E.2 Top-k confidence
Top-$k$ confidence is a likelihood-based OOD detection method that evaluates the probability assigned by the language model to its most confident completions. Given a test input, the model generates $k$ candidate output sequences, each consisting of a sequence of tokens. Let $P_{j}=\prod_{t=1}^{T_{j}}p(y_{t}^{(j)}\mid y_{<t}^{(j)},x)$ denote the joint probability of the $j$-th output sequence, where $x$ is the input, $y_{t}^{(j)}$ is the $t$-th token in the $j$-th output, and $T_{j}$ is the output length. The Top-$k$ confidence score is then defined as:
@@ E.2 Top-k confidence
A lower Top-$k$ confidence score indicates a higher likelihood that the input is out-of-distribution, as the model fails to assign high probability to any of its top candidates.
E.3 Output entropy
Output entropy captures the model’s uncertainty over the space of generated sequences and provides a distributional measure of output dispersion. For the same $k$ candidate sequences used in Top-$k$ confidence, let $P_{j}$ denote the joint probability of the $j$-th output sequence. The normalized distribution over candidates is given by:
@@ E.3 Output entropy
E.4 Last layer embedding
@@
E.5 Best layer embedding
@@ Appendix F GEM Algorithm
1
Require :  GUI agent $\mathcal{F}$, encoder layer $l_{e}$, ID dataset $\mathcal{D}_{\text{ID}}=\{(s_{i},x_{i})\}_{i=1}^{k}$
Ensure :  OOD detection function $f_{\text{OOD}}$
$\mathcal{D}_{\text{embedding}}\leftarrow\emptyset$
for $i\leftarrow 1$ to $k$ do
       $e_{i}\leftarrow l_{e}(s_{i},x_{i})$
       $\mathcal{D}_{\text{embedding}}\leftarrow\mathcal{D}_{\text{embedding}}\cup\{e_{i}\}$
end for
$\mu\leftarrow\frac{1}{k}\sum_{i=1}^{k}e_{i}$
$\mathcal{D}_{\text{distance}}\leftarrow\emptyset$
for $i\leftarrow 1$ to $k$ do
       $d_{i}\leftarrow\lVert e_{i}-\mu\rVert_{2}$
       $\mathcal{D}_{\text{distance}}\leftarrow\mathcal{D}_{\text{distance}}\cup\{d_{i}\}$
end for
for $m\in\{1,\dots,M\}$ do
       Initialize GMM parameters $\{\pi_{j},\mu_{j},\sigma_{j}^{2}\}_{j=1}^{m}$
      repeat
             for $i\leftarrow 1$ to $k$ do
                   for $j\leftarrow 1$ to $m$ do
                         $\gamma_{ij}\leftarrow\frac{\pi_{j}\,\mathcal{N}(d_{i}\mid\mu_{j},\sigma_{j}^{2})}{\sum_{l=1}^{m}\pi_{l}\,\mathcal{N}(d_{i}\mid\mu_{l},\sigma_{l}^{2})}$
                   end for
             end for
            for $j\leftarrow 1$ to $m$ do
                   $\pi_{j}\leftarrow\frac{1}{k}\sum_{i=1}^{k}\gamma_{ij}$
                   $\mu_{j}\leftarrow\frac{\sum_{i=1}^{k}\gamma_{ij}d_{i}}{\sum_{i=1}^{k}\gamma_{ij}}$
                   $\sigma_{j}^{2}\leftarrow\frac{\sum_{i=1}^{k}\gamma_{ij}(d_{i}-\mu_{j})^{2}}{\sum_{i=1}^{k}\gamma_{ij}}$
             end for
      until convergence
      $\log\mathcal{L}_{m}\leftarrow 0$
       for $i\leftarrow 1$ to $k$ do
             $\ell_{i}\leftarrow\sum_{j=1}^{m}\pi_{j}\cdot\mathcal{N}(d_{i}\mid\mu_{j},\sigma_{j}^{2})$
             $\log\mathcal{L}_{m}\leftarrow\log\mathcal{L}_{m}+\log(\ell_{i})$
       end for
      $\text{BIC}(m)\leftarrow-2\log\mathcal{L}_{m}+m\log k$
end for
$m^{*}\leftarrow\arg\min_{m}\text{BIC}(m)$
Fit final GMM with $m^{*}$ components: $\{(\pi_{j},\mu_{j},\sigma_{j})\}_{j=1}^{m^{*}}$
for $j\leftarrow 1$ to $m^{*}$ do
       Define ID interval $I_{j}=[\mu_{j}-3\sigma_{j},\ \mu_{j}+3\sigma_{j}]$
end for
Define $f_{\text{OOD}}(s,x)$ as:
@@
Algorithm 1 GEM Algorithm
Appendix G Ablation Study

### v3

@@ Introduction
Table 2: Comparison of OOD Detection Results Across Different Datasets.
Model Smartphone Computer Web browser
Accuracy(%$\uparrow$ F1 Score(%$\uparrow$ Accuracy(%$\uparrow$ F1 Score(%$\uparrow$ Accuracy(%$\uparrow$ F1 Score(%$\uparrow$
@@
Table 3: The performance of GEM with different encoder structures.
Pilot experiments show that popular OOD detection methods perform suboptimally in the GUI agent domain. However, we observe that the input embeddings generated by GUI agents naturally lend themselves to modeling the ID representation of $\mathcal{D}_{\text{ID}}$.
Given a GUI agent $\mathcal{F}$ and an ID dataset $\mathcal{D}_{\text{ID}}=\{(s_{i},x_{i})\}_{i=1}^{k}$, we first obtain an encoder layer $l_{e}$ from $\mathcal{F}$. The encoder $l_{e}$ maps each input pair $(s_{i},x_{i})$ to an embedding vector $e_{i}\in{R}^{n}$. Thus, we construct an ID embedding dataset $\mathcal{D}_{\text{embedding}}=\{e_{i}\}_{i=1}^{k}$. Each $e_{i}$ is a point in the $n$-dimensional embedding space.
To model this distribution, we first compute the centroid $\mu$ of $\mathcal{D}_{\text{embedding}}$: $\mu=\frac{1}{k}\sum_{i=1}^{k}e_{i}$. Next, we calculate the Euclidean distance between each embedding $e_{i}$ and the centroid $\mu$: $d_{i}=\lVert e_{i}-\mu\rVert_{2},\quad i=1,\ldots,k$, resulting in a distance dataset $\mathcal{D}_{\text{distance}}=\{d_{i}\}_{i=1}^{k}$.
The distribution of $\mathcal{D}_{\text{distance}}$ is typically multi-centered and may contain multiple modes. Therefore, instead of fitting a simple Gaussian or applying heuristic thresholds, we model it using a GMM. Specifically, we assume that the distances are generated from a mixture of $m$ univariate Gaussian components. The GMM models the probability density function as: $p(d)=\sum_{j=1}^{m}\pi_{j}\,\mathcal{N}(d\mid\mu_{j},\sigma_{j}^{2})$, $\pi_{j}$ is the mixing coefficient of the $j$-th component, satisfying $\sum_{j=1}^{m}\pi_{j}=1$ and $\pi_{j}\geq 0$, $\mathcal{N}(d\mid\mu_{j},\sigma_{j}^{2})$ denotes the density of a univariate Gaussian:
M-step update the parameters:
$\pi_{j}^{\text{new}}\!\!=\frac{1}{k}\sum_{i=1}^{k}\gamma_{ij},~\mu_{j}^{\text{new}}\!\!=\frac{\sum_{i=1}^{k}\gamma_{ij}d_{i}}{\sum_{i=1}^{k}\gamma_{ij}},$ (4)
$\sigma_{j}^{2\,\text{new}}\!\!=\frac{\sum_{i=1}^{k}\gamma_{ij}(d_{i}-\mu_{j}^{\text{new}})^{2}}{\sum_{i=1}^{k}\gamma_{ij}}.$ (5)
$\text{BIC}(m)=-2\log\mathcal{L}_{m}+m\log k,$ (6)
The optimal number of components $m^{*}$ is thus selected as $m^{*}=\arg\min_{m}\text{BIC}(m)$. Once the GMM is fitted with $m^{*}$ components, each Gaussian component $\mathcal{N}(\mu_{j},\sigma_{j}^{2})$ provides a center $\mu_{j}$ and a standard deviation $\sigma_{j}$. We define the ID boundary as the interval $[\mu_{j}-n\sigma_{j},\mu_{j}+n\sigma_{j}]$.
$f_{\text{OOD}}(s_{t},x)=\begin{cases}0,&\text{if }\exists j\text{ s.t. }d_{t}\in[\mu_{j}-n\sigma_{j},\mu_{j}+n\sigma_{j}],\\ 1,&\text{otherwise}.\end{cases}$ (7)
Experiments
Experiments Setup
Datasets.
For the ID dataset, we use the AITZ (Zhang et al. 2024b) dataset, which contains GUI agent data covering more than 70 Android app scenarios. For the OOD datasets, we select eight GUI agent datasets spanning three platforms: smartphone, computer, and web browser. The smartphone platform includes AndroidControl (Li et al. 2024), OS-Kairos (Cheng et al. 2025b), Meta-GUI (Sun et al. 2022), and ScreenSpot-Mobile (Li et al. 2025),while the computer platform includes Omniact-Desktop (Kapoor et al. 2024) and ScreenSpot-Desktop, and the web browser platform includes Omniact-Web and ScreenSpot-Web. The four smartphone datasets simulate extrapolation-OOD scenarios, whereas the computer and web browser datasets simulate internalization-OOD scenarios.
@@
Mode SCROLL(%$\uparrow$ PRESS(%$\uparrow$ STOP(%$\uparrow$ CLICK(%$\uparrow$ TYPE(%$\uparrow$ TOTAL(%$\uparrow$
Zero-Shot 19.24 32.05 0.00 30.00 44.59 26.94
SFT 63.11 32.60 65.35 45.14 44.59 49.38
SFT + GEM 63.110.00↑ 35.002.40↑ 73.2028.06↑ 55.4010.26↑ 52.598.00↑ 58.789.40↑
Table 4: Analysis of experiments to use GEM to improve the end-to-end results. We report the single-step action accuracy of GUI agents for different action types and the total.
Implementation.
Popular GUI Agents (Wu et al. 2024; Qin et al. 2025; Xu et al. 2024) are developed based on Qwen2-VL-7B. To simulate the construction process of a popular GUI Agent, we perform SFT on Qwen2-VL-7B using AITZ train dataset, and then evaluate its OOD detection performance on each sample from eight different OOD datasets and AITZ test dataset (ID). We report the accuracy, precision, recall, and F1 score for each dataset.
Baseline.
We use two commonly-used types of OOD detection methods as our experimental baselines: embedding-based methods and uncertainty-based methods. The embedding-based methods include the TV score (Wang et al. 2024b), last layer embedding, and best layer embedding. The uncertainty-based methods include top-k confidence and output entropy.
Main Results
Further analysis
In this section, we present several interesting observations and propose a potential method for using GEM to enhance end-to-end tasks in GUI agents. We also demonstrate the rationality of GEM through ablation studies (see Appendix).
Generalization Experiment
To demonstrate the generalization ability of GEM, we evaluate its performance across nine different GUI agents or MLLMs. As shown in Table 3, GEM consistently achieves high accuracy across five models spanning three platforms. For the other four models where GEM performs suboptimally, the results remain explainable. OS-Atlas-Base-7B and OS-Atlas-Pro-7B were exposed to most of the smartphone datasets used in our experiment during their GUI grounding pretraining phase, and their visual encoders were not frozen, which contributed to their weaker OOD performance on the smartphone platform. Meanwhile, LLaVA-1.5 and BLIP capture visual information at a coarser granularity. This highlights their limitations in handling the complex embedding space specific to GUI agent tasks.
Layer-Level OOD Detection Analysis
We evaluated the effectiveness of representations extracted from each of the twenty-eight layers of Qwen2-VL-7B for OOD detection using an embedding-based approach. As shown in Figure 4, in most datasets, the AUROC first increases as the layer depth increases, then decreases. However, the AUROC rises again in the last two layers. Interestingly, in 4 out of 8 datasets, the best OOD detection results are achieved at the ninth layer. This might be because, as the layers get deeper, the importance of specific task-related features increases, while the contribution of general visual or textual features decreases. Around the ninth layer, the GUI agent likely finds a best balance.
How to use GEM to improve the end-to-end results
GEM not only enhances the security of GUI agents but also improves the end-to-end results of GUI agents by combining it with other modules. For example, we trained Qwen2-VL-7B using the AITZ dataset and conducted tests using a mixed test set of AITZ and OS-Kairos to simulate a real GUI agent testing environment with multiple knowledge sources. When GEM determines that the current sample is an OOD sample, it requests assistance from the cloud-based GPT-4o to execute tasks. We report the step-wise success rates for different types of GUI agents as well as the total success rate. The experimental results are shown in Table 4, indicating that GEM can enhance the step-wise success rate of GUI agents by 9.4%, demonstrating the potential of GEM to improve the end-to-end results.
Time Cost of GEM
Although GEM calculates the distances in the embedding space multiple times, it does not actually incur a burden on the time overhead. We conducted experiments on the time cost of GEM, and the experimental results are shown in Table 5. The experiments indicate that GEM only increases training time by 4.9% and testing time by 6.5%.
Training time Testing time
Original time 3.63 h 0.77 s
Time with GEM 0.18 h 0.05 s
Additional time (%) 4.9 6.5
Table 5: Time Comparison of GUI Agent w/ and w/o GEM.
Conclusion
We present the first systematic analysis of OOD detection for GUI agents and show that traditional OOD detection methods perform suboptimally in this domain. Then we propose GEM, a method that uses input embedding information to fit a GMM to detect OOD samples for GUI agents. Experiments show that GEM achieves an average accuracy improvement of 23.70% over the best-performing baseline while only increasing training time by 4.9% and testing time by 6.5%. And we present several interesting observations and propose a potential method for using GEM to enhance end-to-end tasks in GUI agents.
Appendix
Ethics Statement
@@ Conclusion
Limitations
@@
Expanded Analysis
The detailed analysis of clustering phenomenon
@@
Figure 5: GUI agent performance across different cluster distances.
Why does the TV score fail in OOD detection for GUI agents?
GUI agents operate on terminal devices with very similar thinking logic. The TV score assesses the model’s thought process based on the changes in the embeddings of adjacent layers, but it cannot distinguish whether the current scene is OOD or not. As shown in Figure 6, the TV score is indistinguishable between OOD samples and ID samples.
Datasets Details
@@
AITZ is the first dataset to annotate Chain-of-Action-Thought (CoAT) for GUI agents, which considers descriptions of prior actions and the current screen, the actions that should be taken, and reasoning about the potential consequences of the selected actions. It contains 18,643 screen-action pairs with CoAT annotations.
AndroidControl is a dataset covering 833 distinct Android applications and containing 15,283 everyday tasks performed within these apps. This dataset is currently the most diverse benchmark for computer-based control, providing a foundation for in-depth model performance analysis.
OS-Kairos is a dataset containing data from 12 different Chinese and English mobile applications across 12 topic categories. The tasks in this dataset were automatically collected by an agent capability probing framework and later curated by humans. It covers a wide range of scenarios encountered in everyday life.
Meta-GUI is a dataset containing dialogues and GUI interaction trajectories. It covers six high-frequency daily-life topics across 11 mobile applications.
Omniact is a dataset consisting of 9.8K pairs of screen images and natural language task descriptions, covering a variety of operating systems and web applications. The dataset includes tasks of varying difficulty levels, ranging from simple single-step actions to complex multi-step operations.
ScreenSpot is a dataset containing high-resolution real screenshots and expert annotations from various domains. This dataset covers data from multiple platforms, including Mac, Windows, Android, iPhone, and web, making it the most platform-diverse GUI dataset.
Figure 6: The distinguishing effect of the TV score on AndroidControl and AITZ.
Main Experiment Details
@@
Evaluation Metrics Details. After performing SFT on Qwen2-VL-7B using the AITZ training set, we fit a distance-based GMM on the same training set following the GEM method. Classification boundaries are defined using three standard deviations per cluster. The accuracy, precision, recall, and F1 scores reported in main results are obtained by treating each dataset as the OOD samples and the AITZ test set as the ID samples. The prompt of GEM and all baselines is shown in Figure 7.
@@
Top-k confidence
[…]
@@
[…]
Output entropy
[…]
@@
Last layer embedding
@@
Best layer embedding
@@
Algorithm 1 GEM Algorithm
1:GUI agent $\mathcal{F}$, encoder layer $l_{e}$, ID dataset $\mathcal{D}_{\text{ID}}=\{(s_{i},x_{i})\}_{i=1}^{k}$
2:OOD detection function $f_{\text{OOD}}$
[…]
4:for $i\leftarrow 1$ to $k$ do
5:  $e_{i}\leftarrow l_{e}(s_{i},x_{i})$
6:  $\mathcal{D}_{\text{embedding}}\leftarrow\mathcal{D}_{\text{embedding}}\cup\{e_{i}\}$
7:end for
[…]
10:for $i\leftarrow 1$ to $k$ do
11:  $d_{i}\leftarrow\lVert e_{i}-\mu\rVert_{2}$
12:  $\mathcal{D}_{\text{distance}}\leftarrow\mathcal{D}_{\text{distance}}\cup\{d_{i}\}$
13:end for
14:for $m\in\{1,\dots,M\}$ do
15:  Initialize GMM parameters $\{\pi_{j},\mu_{j},\sigma_{j}^{2}\}_{j=1}^{m}$
16:  repeat
17:   for $i\leftarrow 1$ to $k$ do
18:     for $j\leftarrow 1$ to $m$ do
19:      $\gamma_{ij}\leftarrow\frac{\pi_{j}\,\mathcal{N}(d_{i}\mid\mu_{j},\sigma_{j}^{2})}{\sum_{l=1}^{m}\pi_{l}\,\mathcal{N}(d_{i}\mid\mu_{l},\sigma_{l}^{2})}$
20:     end for
21:   end for
22:   for $j\leftarrow 1$ to $m$ do
23:     $\pi_{j}\leftarrow\frac{1}{k}\sum_{i=1}^{k}\gamma_{ij}$
24:     $\mu_{j}\leftarrow\frac{\sum_{i=1}^{k}\gamma_{ij}d_{i}}{\sum_{i=1}^{k}\gamma_{ij}}$
25:     $\sigma_{j}^{2}\leftarrow\frac{\sum_{i=1}^{k}\gamma_{ij}(d_{i}-\mu_{j})^{2}}{\sum_{i=1}^{k}\gamma_{ij}}$
26:   end for
27:  until convergence
28:  $\log\mathcal{L}_{m}\leftarrow 0$
29:  for $i\leftarrow 1$ to $k$ do
30:   $\ell_{i}\leftarrow\sum_{j=1}^{m}\pi_{j}\cdot\mathcal{N}(d_{i}\mid\mu_{j},\sigma_{j}^{2})$
31:   $\log\mathcal{L}_{m}\leftarrow\log\mathcal{L}_{m}+\log(\ell_{i})$
32:  end for
33:  $\text{BIC}(m)\leftarrow-2\log\mathcal{L}_{m}+m\log k$
34:end for
[…]
36:Fit final GMM with $m^{*}$ components: $\{(\pi_{j},\mu_{j},\sigma_{j})\}_{j=1}^{m^{*}}$
37:for $j\leftarrow 1$ to $m^{*}$ do
38:  Define ID interval $I_{j}=[\mu_{j}-3\sigma_{j},\ \mu_{j}+3\sigma_{j}]$
39:end for
40:Define $f_{\text{OOD}}(s,x)$ as:
@@
Ablation Study
@@
Why Qwen2-VL-2B Performs Better than Qwen2.5-VL-3B for GEM?
The performance discrepancy between Qwen2-VL-2B and Qwen2.5-VL-3B in our encoder comparison experiments stems from fundamental architectural differences in their vision encoders. While both models belong to the Qwen-VL family, their encoders are not identical:
Qwen2-VL-2B employs a standard ViT (Vision Transformer) architecture.
Qwen2.5-VL-3B uses a custom window-based attention encoder, where images are split into non-overlapping windows (112×112, equivalent to 8×8 patches). Each window computes self-attention locally, with no cross-window interaction initially, followed by a late-stage full-attention block.
Since GEM relies solely on the encoder component, the model’s total parameter count (e.g., 2B vs. 3B) does not directly predict its effectiveness in our framework. The superior performance of Qwen2-VL-2B suggests that ViT’s global attention mechanism is better suited for GEM’s objectives compared to Qwen2.5-VL-3B’s hybrid window/full-attention design.